A motorcycle has a constant acceleration of 4.45 m/s2. Both the velocity and acceleration of the motorcycle point in the same direction. How much time is required for the motorcycle to change its speed from (a)31.0 to 41.0 m/s, and (b)61.0 to 71.0 m/s?
Answer #1

a)V=U+at where V is final velocity and U is initial velocity

T=(V-U)/a=(41-31)/4.45=2.247sec=2.25 sec

b)T=(V-U)/a=(71-61)/4.45=2.247sec=2.25 sec